Find the magnitude and direction of the resultant of the following vectors using Laws of Sines and Cosines:10 N, E50°N and 8 N, W10°S.


1
Expert's answer
2021-04-13T06:29:14-0400

Explanations & Calculations


  • Refer to the sketch attached.


  • Two forces can be rearranged in a vector triangle as shown where laws of sines & cosines can be applied then.
  • Due to the geometry the angle BOA = 40 degrees
  • And the angle OBA can be found to be "\small (190-\theta)"
  • Then by cosine rule, the magnitude of the resultant can be found

"\qquad\qquad
\begin{aligned}
\small \cos40&=\small \frac{OB^2+OA^2-AB^2 }{2OB\times OA}\\
\small 0.7660&=\small \frac{8^2+10^2-R^2}{2\times 8\times 10}\\
\small R&=\small \bold{6.437\,N}
\end{aligned}"

  • If it makes an angle "\small \theta" with the north, then using sine rule,

"\qquad\qquad
\begin{aligned}
\small \frac{10}{\sin(190-\theta)}&=\small \frac{R}{\sin40}\\
\small\sin(190-\theta)&=\small 0.99858\\
\small 190-\theta &=\small 86.946\\
\small \theta&=\small \bold{103.05^0[N\,of\,E]}
\end{aligned}"